OpenStudy (anonymous):

Can anyone figure out what x and y are in this problem using th method: Elimination.? 7x - 3y = 20 5x + 3y = 16 Thanks.(:

OpenStudy (anonymous):

is this algebra 1??

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

by sum : 7x - 3y = 20 5x + 3y = 16 -------------- 12x+0=36 x=36/12=3

OpenStudy (anonymous):

I know you subtract and end up with 12x = 36 or am I wrong?

OpenStudy (anonymous):

ok well, um thats it @baziham lol :)

OpenStudy (anonymous):

and by equation 1: 7x - 3y = 20 7(3) - 3y=20 21 - 20 = 3y 1= 3y y= 1/3

OpenStudy (anonymous):

Whats x then?

OpenStudy (anonymous):

or is there no x? :o

OpenStudy (anonymous):

x=3 i write it at the end of first post ok?

OpenStudy (phi):

I know you subtract and end up with 12x = 36 or am I wrong? almost. you add. you can do it like baz showed,(which is how I would do it) or like this, all on one line 7x +5x - 3y +3y = 20+16 simplify and you get 12x=36 and x= 36/12 or 3

OpenStudy (phi):

now use x=3 in either equation and solve for y

OpenStudy (anonymous):

Oh, I gotcha. And I ment to say add. Lol.

OpenStudy (phi):

and the reason we add is because we noticed that -3y + 3y is zero, making the y's disappear.

OpenStudy (anonymous):

Oh! Thankyou! You all are such a big help.! I was so confused.